# Ultrafast Solvent Dynamics Drives the Formation of the Hydrated Electron in Photoexcited Water

### v2

@@ 2.3 H Atom Transfer Mechanism
The most probable configuration, with $CN_{H}=1$ and $CN_{O}=2$ (Figure 3A), corresponds to the initial state of the simulation, where the hydrogen atom is still part of a neutral water molecule. Upon photoexcitation, we observe a fast dissociation of this hydrogen atom, accompanied by a sharp drop in the $S_{1}\to S_{0}$ energy gap from approximately 8 eV to below 0.2 eV, leading to a non-radiative transition (red points in the Figure 3). In some trajectories, the dissociating hydrogen atom moves into an empty cavity (Figure 3B), where its nearest oxygen is part of a hydroxyl radical ($\mathrm{HO}^{\bullet}$), resulting in $CN_{O}=1$.
Another frequently observed pathway is the formation of a species in which the dissociating hydrogen atom binds to a water molecule, forming the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$), characterized by $CN_{H}\sim$1 and $CN_{O}\sim$3 (Figure 3C). We confirmed that this species is indeed $\mathrm{H_{3}O}^{\bullet}$ and not the conventional $\mathrm{H_{3}O}^{+}$ ion, since the electron remains associated with it. While the species $\mathrm{H_{3}O}^{\bullet}$ has been predicted by various theoretical models28, 56, 57, 70, our work is the first to demonstrate its formation following excitation by simulating the entire process from photoabsorption to relaxation in liquid water. According to our simulations, the radical $\mathrm{H_{3}O}^{\bullet}$ appears only transiently. It eventually undergoes further dissociation, where the hydrogen atom carrying the electron migrates into a region surrounded by water molecules. In these configurations, an OH bond from a nearby water molecule points toward the hydrogen atom (see Figure 3D) creating an unstable geometry ultimately leading to non-radiative decay to the ground-state. To the best of our knowledge the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$) has not been observed experimentally through photoionization/photoabsorption of water perhaps in part due to the fact that it is formed fleetingly as a rare transition state-like motif.
@@ 2.4 Proton Coupled Electron Transfer Mechanism
Experimental studies have reported initial gyration radius of around 20-40 Å34, 67 immediately after photoabsorption. Although our method captures the same qualitative behavior, direct comparison with experiment is limited by the size of the simulation box (12.42 Å), which restricts the maximum observable extent of delocalization71. The final gyration radius we observe in the excited state is 2.7 Å, slightly larger than the experimentally reported value of 2.4 Å for the hydrated electron in the ground state72. Note that the decrease of the gyration radius from 6 to 2.4 Å incurs a decrease in the energy gap. Once the electron has localized, the system undergoes a non-radiative transition to the ground state, as indicated by the red points in Figure 4A.
@@
As mentioned in the introduction, generating a free electron in bulk water using photon energies within the first absorption band requires some degree of nuclear rearrangement to release the electron from the molecule17. Earlier, we showed that this release occurs simultaneously with the dissociation of the OH bond in the excited water molecule, leading to the formation of a free electron, a hydroxyl radical ($\mathrm{HO}^{\bullet}$), and a proton ($\mathrm{H}^{+}$or $\mathrm{H_{3}O}^{+}$). We now move on to elucidating that the localization of the electron and the formation of the other species requires a collective process involving reorganization of the hydrogen bond network of the water molecules.
@@
In Figure 4B we illustrate the coupling between the gyration radius against the angle formed between the vector connecting the oxygen and the center of the electron and the OH vector of the water molecule. For this analysis we only include the nearby water molecules involved in the first solvation shell of the electron, (see Analysis Protocols section for details). When the electron is delocalized, this angle spans a broad range from 0 to 120, reflecting the absence of a well-defined solvation shell. As the system relaxes, the angle narrows to below 30, consistent with the orientational preferences seen in hydrated electron solvation shells58, 71, 73. These results indicate that the structure of the hydrated electron forms within the excited-state itself, before the non-radiative decay consistent with recent experiments using sub-two-cycle visible-to-shortwave infrared pump–probe spectroscopy74. Analysis of the radial distribution function in the low gyration radius regime on the excited state (see Figure S9 in the Supplementary Information), where the hydrated electron is effectively localized, reveals that the first solvation shell comprises between 4-5 water molecules. This is consistent with previous findings for the hydrated electron in the ground electronic state45, 51, 25.
@@
Figure 4: Coupled Electron Proton Transfer Mechanism. Panel A: Probability Density Function of the gyration radius and the energy gap for all the trajectories. Red circles represent the conformations at the crossing point. Insets show the electronic spin density for representative conformations of the system at two different times during the excited state dynamics and at different iso-values. Panel B: Probability Density Function of the gyration radius and the angle formed between the vector pointing from the Oxygen to the electron center and the vector of the O-H of the waters in the first solvation shell. Red circles represent the conformations at the crossing point. Insets show schematic representations of the water at the two gyration radius along with the vectors considered for the analysis.
@@
2.5 Coupled Translation and Rotational Solvent Dynamics Stabilizes Water-Mediated Photoproducts
@@
In addition to these rotational dynamics, Figure 5B shows that localization also involves collective translational motion. Water molecules in the first solvation shell move on average by approximately 1.2 Å relative to their final positions, a displacement slightly larger than that of the more distant water molecules. Overall, our results show that both rotational and translational rearrangements of specific water molecules are necessary for the localization of the hydrated electron. These collective motions also cause a notable distortion of the hydrogen bond network, consistent with experimental findings from ionization studies in liquid water35.
@@
As alluded to earlier, the creation of the excess electron after photoexcitation also leads to the formation of a hydronium ion ($\mathrm{H_{3}O}^{+}$) and a hydroxyl radical ($\mathrm{HO}^{\bullet}$). To investigate the behavior of these species during the excited-state dynamics, we present in Figure 5C the coupling between the gyration radius of the electron and the distance between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H_{3}O}^{+}$. Interestingly, our analysis reveals three distinct minima in the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance, particularly in the region of low gyration radius where the electron becomes localized. These minima occur at $\sim$2.6, 4.3, and 6.2 Å. Our simulations are in excellent agreement with very recent time-resolved electron diffraction experiments following the ionization of liquid water, where the corresponding minima were found at 2.4, 3.4, and 5.8 Å35. These three minima correspond to situations involving the initial formation of contact ion-radical pair and then the separation mediated by the solvent. Although previous theoretical studies have identified the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair by simulating ionized water with the electron removed75, 76, 77, our findings show for the first time, the creation of these correlated structures upon photoexcitation starting from the neutral bulk water.
@@
The creation of ion-radical pairs at different distances has important implications on the optical properties. To probe this, we show in Figure 5D the energy gap as a function of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance. For the contact ion-radical pair at 2.6 Å the average energy gap is $\sim$3 eV on average. As one moves towards the creation of the solvent-separated ion-radical pairs, the energy gap drops to $\sim$2.0 eV. This trend is also consistent with experimental observations, which demonstrated that the formation and spatial separation of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair occur before the internal conversion associated with the $p\to s$ transition of the hydrated electron35.
@@ 2.6 Spontaneous Emission from Hydrated Excess Electrons
An important and still not fully understood property of the hydrated electron in liquid water is its photon emission following excitation. Understanding this emission can provide valuable insight into the nature of the excited state and the relaxation dynamics of the system. Although experimental studies are relatively scarce, Tauber and Mathies have explored this optical behavior through fluorescence and Raman spectroscopies58. Motivated by their findings, we now focus on analyzing the emission spectrum obtained from our simulations and examine how it depends on the gyration radius of the hydrated electron.
@@
Figure 6A shows the experimental emission spectrum along with our calculated $S_{1}\to S_{0}$ energy gaps, evaluated for configurations with both large and small electron gyration radius. As the excess electron becomes more localized, transitioning from a larger to a smaller gyration radius, the energy gap decreases from an average of 2.4 to 1.3 eV (516 to 954 nm). Interestingly, the experimental emission spectrum is in excellent agreement with our predictions in both peak position and spectral width, especially for configurations with smaller gyration radius. This observation supports previous interpretations from experimental58 and theoretical studies78, 79, suggesting that emission occurs from a broad ensemble of transient excited state geometries rather than from a single well-defined minimum. It is important to highlight that our theoretical spectrum reflects the recombination of the electron, since we describe the final electronic ground state as a closed shell (see Computational Methods). In contrast, the experimental measurements capture emission from the excited state of the hydrated electron without identifying the final state58, which could involve either recombination or hydrated electron in the ground state. Therefore, while the agreement between our results and the experimental data is encouraging, it should be considered preliminary. A more detailed investigation of these optical properties is necessary and lies beyond the scope of this study. However, our findings provide a valuable framework that could guide future experimental and theoretical efforts aimed at confirming and better understanding the emission behavior of the hydrated electron.
@@
In order to ensure that these features are not sensitive to the choice of our ROKS method used for the simulation of the excited state, we re-examined the energy gap along the time for one representative trajectory undergoing the PCET mechanism using TDDFT approach. Figure 6B shows that the two approaches are in excellent agreement, with an average deviation of only 0.2 eV. We also examined the particle orbital obtained from TDDFT and compared it with the spin density from ROKS. As shown in the inset of Figure 6B, both orbitals exhibit a similar spatial distribution and involving the same water molecules in the first solvation shell of the electron. We also computed the emission spectrum for the last 300 fs of the same trajectory using the energy gaps and oscillator strengths obtained from TDDFT (Figure 6C). While the spectrum is not fully converged due to limited sampling, the resulting peak at 1.7 eV closely matches ROKS calculated and experimental emission peak.
@@
Figure 6: Photon emission properties of the hydrated electron. Panel A: Histograms of the energy gap using all the frames of all the trajectories that undergo the PCET mechanism. Light blue bar is for all the frames with a gyration radius lower than 4 Å. Light purple bar is for all the frames with a gyration radius higher than 4 Å. Green and red solid lines are the experimental and fitted fluorescence spectra of the hydrated electron. Experimental data were digitalized from Ref. 58. Panel B: Time dependent evolution of the energy gap for a characteristic trajectory using ROKS method (green) and TDDFT (purple). The insets show the spin density and the particle orbital of the hydrated electron at 250 fs (denoted as a red point) for both ROKS and TDDFT methods, respectively. Panel C: Fluorescence spectra using the oscillator strength and energy gap for the last 300 fs for this trajectory using TDDFT.

### v3

@@ 2.3 H Atom Transfer Mechanism
The most probable configuration, with $CN_{H}=1$ and $CN_{O}=2$ (Figure 3A), corresponds to the initial state of the simulation, where the hydrogen atom is still part of a neutral water molecule. Upon photoexcitation, we observe a fast dissociation of this hydrogen atom, accompanied by a sharp drop in the $S_{1}\to S_{0}$ energy gap from approximately 8 eV to below 0.2 eV, leading to a non-radiative transition (as indicated by all the red points in the Figure 3). In some trajectories, the dissociating hydrogen atom moves into an empty cavity (Figure 3B), where its nearest oxygen is part of a hydroxyl radical ($\mathrm{HO}^{\bullet}$), resulting in $CN_{O}=1$ and leading to a non-radiative crossing to the ground state.
Another frequently observed pathway is the formation of a species in which the dissociating hydrogen atom binds to a water molecule, forming the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$), characterized by $CN_{H}\sim$1 and $CN_{O}\sim$3 (Figure 3C). We confirmed that this species is indeed $\mathrm{H_{3}O}^{\bullet}$ and not the conventional $\mathrm{H_{3}O}^{+}$ ion, since the electron remains associated with it. While the species $\mathrm{H_{3}O}^{\bullet}$ has been predicted by various theoretical models42, 71, 72, 89, our work is the first to demonstrate its formation following excitation by simulating the entire process from photoabsorption to relaxation in liquid water. According to our simulations, the radical $\mathrm{H_{3}O}^{\bullet}$ appears only transiently. It eventually undergoes further dissociation, where the hydrogen atom carrying the electron migrates into a region surrounded by water molecules. In these configurations, an OH bond from a nearby water molecule points toward the hydrogen atom (see Figure 3D) creating an unstable geometry ultimately leading to non-radiative decay to the ground-state. To the best of our knowledge the radical hydronium ($\mathrm{H_{3}O}^{\bullet}$) has not been observed experimentally through photoionization/photoabsorption of water perhaps in part due to the fact that it is formed fleetingly as a rare transition state-like motif.
@@
An interesting observation is that approximately 20% of the trajectories follow the mechanism involving the formation of a hydrogen atom in an empty cavity (Figure 3B) with a lifetime of 13 fs, while the remaining 33% correspond to the alternative HAT pathway (Figures 3C and 3D) exhibiting a longer lifetime of 34 fs. These results correlate well with the defect analysis presented in Figure 1C. When the dissociating water molecule has a missing hydrogen-bond donor (AD or AAD defect), there is sufficient space available to allow an ultrafast hydrogen atom dissociation into a cavity ($\sim$13 fs), in excellent agreement with values reported for a hydrogen-bonded water dimer in vacuum88, 54. In contrast, when the dissociating molecule lacks a hydrogen-bond acceptor (ADD defect), the surrounding hydrogen-bond network in the condensed phase restricts this motion, delaying the non-radiative decay and promoting the formation of the transient $\mathrm{H_{3}O}^{\bullet}$ species.
@@ 2.4 Proton Coupled Electron Transfer Mechanism
Experimental studies have reported initial gyration radius of around 20-40 Å85, 75 immediately after photoabsorption. Although our method captures the same qualitative behavior, direct comparison with experiment is limited by the size of the simulation box (12.42 Å), which restricts the maximum observable extent of delocalization10. The final gyration radius we observe in the excited state is 2.7 Å, slightly larger than the experimentally reported value of 2.4 Å for the hydrated electron in the ground state21. Note that the decrease of the gyration radius from 6 to 2.4 Å incurs a decrease in the energy gap. Once the electron has localized, the system undergoes a non-radiative transition to the ground state, as indicated by the red points in Figure 4A.
@@
As mentioned in the introduction, generating a free electron in bulk water using photon energies within the first absorption band requires some degree of nuclear rearrangement to release the electron from the molecule27. Earlier, we showed that this release occurs simultaneously with the dissociation of the OH bond in the excited water molecule, leading to the formation of a free electron, a hydroxyl radical ($\mathrm{HO}^{\bullet}$), and a proton ($\mathrm{H}^{+}$or $\mathrm{H_{3}O}^{+}$). We now move on to elucidating that the localization of the electron and the formation of the other species requires a collective process involving reorganization of the hydrogen bond network of the water molecules.
@@
In Figure 4B we illustrate the coupling between the gyration radius against the angle formed between the vector connecting the oxygen and the center of the electron and the OH vector of the water molecule. For this analysis we only include the nearby water molecules involved in the first solvation shell of the electron, (see Analysis Protocols section for details). When the electron is delocalized, this angle spans a broad range from 0 to 120, reflecting the absence of a well-defined solvation shell. As the system relaxes, the angle narrows to below 30, consistent with the orientational preferences seen in hydrated electron solvation shells102, 10, 48. These results indicate that the structure of the hydrated electron forms within the excited-state itself, before the non-radiative decay consistent with recent experiments using sub-two-cycle visible-to-shortwave infrared pump–probe spectroscopy66. Analysis of the radial distribution function in the low gyration radius regime on the excited state (see Figure S9 in the Supplementary Information), where the hydrated electron is effectively localized, reveals that the first solvation shell comprises between 4-5 water molecules. This is consistent with previous findings for the hydrated electron in the ground electronic state58, 17, 76.
@@
Figure 4: Proton Coupled Electron Transfer Mechanism. Panel A: Probability Density Function of the gyration radius and the energy gap for all the trajectories. Red circles represent the conformations at the crossing point. Insets show the electronic spin density for representative conformations of the system at two different times during the excited state dynamics and at different iso-values. Panel B: Probability Density Function of the gyration radius and the angle formed between the vector pointing from the Oxygen to the electron center and the vector of the O-H of the waters in the first solvation shell. Red circles represent the conformations at the crossing point. Insets show schematic representations of the water at the two gyration radius along with the vectors considered for the analysis.
@@
2.5 Coupled Translational and Rotational Solvent Dynamics Stabilizes Water-Mediated Photoproducts
@@
In addition to these rotational dynamics, Figure 5B shows that localization also involves collective translational motion. Water molecules in the first solvation shell move on average by approximately 1.2 Å relative to their final positions, a displacement slightly larger than that of the more distant water molecules. Overall, our results show that both rotational and translational rearrangements of specific water molecules are necessary for the localization of the hydrated electron. These collective motions also cause a notable distortion of the hydrogen bond network, consistent with experimental findings from ionization studies in liquid water62. A similar reorganizational response has been reported in ab initio molecular dynamics simulations of carbon dioxide reduction in water82, 73, where the breaking of hydrogen bonds in the second solvation shell triggers the stabilization of the excess electron on the carbon dioxide molecule. This process also involves coupled translational and rotational motions of water molecules, closely paralleling the behavior observed in our simulations.
@@
As alluded to earlier, the creation of the excess electron after photoexcitation also leads to the formation of a hydronium ion ($\mathrm{H_{3}O}^{+}$) and a hydroxyl radical ($\mathrm{HO}^{\bullet}$). To investigate the behavior of these species during the excited-state dynamics, we present in Figure 5C the coupling between the gyration radius of the electron and the distance between the $\mathrm{HO}^{\bullet}$ and the $\mathrm{H_{3}O}^{+}$. Interestingly, our analysis reveals three distinct minima in the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance, particularly in the region of low gyration radius where the electron becomes localized. These minima occur at $\sim$2.6, 4.3, and 6.2 Å. Our simulations are in excellent agreement with very recent time-resolved electron diffraction experiments following the ionization of liquid water, where the corresponding minima were found at 2.4, 3.4, and 5.8 Å62. These three minima correspond to situations involving the initial formation of contact ion-radical pair and then the separation mediated by the solvent. Although previous theoretical studies have identified the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair by simulating ionized water with the electron removed67, 78, 47, our findings show for the first time, the creation of these correlated structures upon photoexcitation starting from the neutral bulk water.
@@
The creation of ion-radical pairs at different distances has important implications on the optical properties. To probe this, we show in Figure 5D the energy gap as a function of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) distance. For the contact ion-radical pair at 2.6 Å the average energy gap is $\sim$3 eV on average. As one moves towards the creation of the solvent-separated ion-radical pairs, the energy gap drops to $\sim$2.0 eV. This trend is also consistent with experimental observations, which demonstrated that the formation and spatial separation of the $\mathrm{HO}^{\bullet}$($\mathrm{H_{3}O}^{+}$) pair occur before the internal conversion associated with the $p\to s$ transition of the hydrated electron62.
@@ 2.6 Spontaneous Emission from Hydrated Excess Electrons
An important and still not fully understood property of the hydrated electron in liquid water is its photon emission following excitation. Understanding this emission can provide valuable insight into the nature of the excited state and the relaxation dynamics of the system. Although experimental studies are relatively scarce, Tauber and Mathies have explored this optical behavior through fluorescence and Raman spectroscopies102. Motivated by their findings, we now focus on analyzing the emission spectrum obtained from our simulations and examine how it depends on the gyration radius of the hydrated electron.
@@
Figure 6A shows the experimental emission spectrum along with our calculated $S_{1}\to S_{0}$ energy gaps, evaluated for configurations with both large and small electron gyration radius. As the excess electron becomes more localized, transitioning from a larger to a smaller gyration radius, the energy gap decreases from an average of 2.4 to 1.3 eV (516 to 954 nm). Interestingly, the experimental emission spectrum is in excellent agreement with our predictions in both peak position and spectral width, especially for configurations with smaller gyration radius. This observation supports previous interpretations from experimental102 and theoretical studies87, 29, suggesting that emission occurs from a broad ensemble of transient excited state geometries rather than from a single well-defined minimum. It is important to highlight that our theoretical spectrum reflects the recombination of the electron, since we describe the final electronic ground state as a closed shell (see Computational Methods). In contrast, the experimental measurements capture emission from the excited state of the hydrated electron without identifying the final state102, which could involve either recombination or hydrated electron in the ground state. Therefore, while the agreement between our results and the experimental data is encouraging, it should be considered preliminary. A more detailed investigation of these optical properties is necessary and lies beyond the scope of this study. However, our findings provide a valuable framework that could guide future experimental and theoretical efforts aimed at confirming and better understanding the emission behavior of the hydrated electron.
@@
In order to ensure that these features are not sensitive to the choice of our ROKS method used for the simulation of the excited state, we computed the emission spectrum for the last 300 fs of a representative PCET trajectory using the energy gaps and oscillator strengths obtained from TDDFT (Figure 6B). While the spectrum is not fully converged due to limited sampling, the resulting peak at 1.7 eV closely matches ROKS calculated and experimental emission peak. Additional validations for both mechanisms are provided in the next section.
@@
Figure 6: Photon emission properties of the hydrated electron. Panel A: Histograms of the energy gap using all the frames of all the trajectories that undergo the PCET mechanism. Light blue bar is for all the frames with a gyration radius lower than 4 Å. Light purple bar is for all the frames with a gyration radius higher than 4 Å. Green and red solid lines are the experimental and fitted fluorescence spectra of the hydrated electron. Experimental data were digitalized from Ref. 102. Panel B: Fluorescence spectra using the oscillator strength and energy gap for the last 300 fs for this trajectory using TDDFT.
3 Evaluating the Accuracy of ROKS for the Photochemistry of Liquid Water
In this work, we employed the ROKS approach to investigate the photochemistry of liquid water. While our results show excellent qualitative and quantitative agreement with a broad range of experimental observations, it is important to assess the quality of the electronic structure method and to evaluate its performance relative to more extensively tested theoretical approaches in photochemistry.
In order to first assess the robustness of ROKS with respect to the choice of exchange–correlation functional and basis set, we performed additional calculations using the CAM-B3LYP functional118 and the TZV2P basis set107. The results, presented in section S9 in the Supplementary Information, demonstrate that the ROKS method provides consistent energy gaps and reproduces the same behavior for both photochemical pathways (HAT and PCET), confirming its reliability across different electronic structure settings.
We further compared our results with the TDDFT method, which has been widely applied to the study of excited states and in non-adiabatic dynamics. In addition, for a hydrogen-bonded water dimer in vacuum, we benchmarked our approach against both multiconfigurational methods and TDDFT. The main results are summarized in Figure 7.
Figure 7: Reliability of the ROKS approach for describing the photochemistry of water. Panel A: Evolution of the energy gap for a representative HAT trajectory obtained with ROKS (green) and TDDFT (purple). Panel B: Evolution of the energy gap for a representative PCET trajectory obtained with ROKS (green) and TDDFT (purple). Panel C: Spin density from ROKS and LUMO orbital from TDDFT at 250 fs of the PCET trajectory, illustrating the localization of the hydrated electron. Panel D: Minimum Energy Path (MEP) for a water dimer in vacuum exhibiting the HAT mechanism, obtained with CASPT2. The inset shows the geometry at the conical intersection between $S_{1}$ (purple) and $S_{0}$ (red). Data reproduced from ref. 88. Panel E: Linear Interpolated Path (LIP) for the same water dimer in vacuum exhibiting the HAT mechanism, obtained with ROKS-CAM-B3LYP. The inset shows the geometry at the $S_{1}/S_{0}$ conical intersection. Panel F: Linear Interpolated Path (LIP) for the water dimer exhibiting the HAT mechanism, obtained with TDDFT-CAM-B3LYP. The inset shows the geometry at the $S_{1}/S_{0}$ conical intersection.
Panels 7A and 7B show the evolution of the energy gap between the first excited state ($S_{1}$) and the ground state ($S_{0}$) for two representative trajectories (sampled from the ROKS dynamics), corresponding to the HAT and PCET mechanisms, obtained using both ROKS and TDDFT with the same functional and basis set employed in this work (see Computational Methods section). At the initial step, both methods yield nearly identical energy gaps, as further confirmed by a broader validation across all initial conditions (see Figure S1 in the Supplementary Information).
The main difference arises in the HAT mechanism. At the end of the simulation, ROKS predicts a crossing point with an energy gap smaller than $0.2$ eV, whereas TDDFT yields a value near $1.5$ eV for the same geometry. Despite this quantitative difference, both methods describe the same physical process: water dissociation leading to the formation of $\mathrm{HO}^{\bullet}$ and $\mathrm{H}^{\bullet}$. In contrast, the PCET mechanism shows excellent agreement between ROKS and TDDFT. In this case, both methods predict the dissociation of a water molecule into $\mathrm{HO}^{\bullet}$ and $\mathrm{H_{3}O}^{+}$ within the first 20 fs, with comparable energy gap profiles. Moreover, TDDFT also reproduces the localization of the hydrated electron, consistent with the ROKS results, as shown in Figure 7C, which compares the ROKS spin density and TDDFT LUMO at 250 fs.
To further understand the discrepancy between ROKS and TDDFT observed for the HAT mechanism, we examined a model system consisting of a hydrogen-bonded water dimer in vacuum 88, 54. Segarra et al. computed the Minimum Energy Path (MEP) for the photodissociation of one water molecule from the Franck–Condon region to the $S_{1}/S_{0}$ conical intersection using the multiconfigurational CASPT2 method (see Figure 7D). Using both the ground-state and conical intersection optimized geometries at this high level of theory, we computed the potential energy surfaces along the Linear Interpolated Path (LIP) with ROKS (Figure 7E) and TDDFT (Figure 7F) employing the CAM-B3LYP functional. It is worth emphasizing that the numerical values along horizontal axes for the MEP and LIP coordinates do not correspond to identical nuclear configurations, except for the initial point shown in Figure 7D-F. We thus focus on the overall trends of the ground and excited state potential energy surfaces.
ROKS predicts a dissociative pathway along the $S_{1}$ surface, similar in slope to CASPT2. The main difference lies in the position of the crossing point: at a $\mathrm{HO}^{\bullet}$$\mathrm{H}^{\bullet}$ distance of approximately 2 Å for ROKS, compared with 2.8 Å for CASPT2. Although an O-H distance of 2 Å could be considered as dissociated geometry, the topology of the conical intersection differs between the two methods. Despite this difference, ROKS reproduces the same essential physical process, namely the hydrogen atom dissociation. In contrast, TDDFT predicts a local minimum on the $S_{1}$ surface at around 2 Å with an energy gap of $\sim 1.5$ eV, which is consistent with the differences seen in the condensed phase simulations for the HAT mechanism (Figure 7A). Moreover, TDDFT exhibits a $\sim 1$ eV barrier from this minimum to the conical intersection, which could significantly delay, or even prevent, reaching the intersection within the timescales ($\sim 10$ fs) reported in previous studies 88, 54.
Taken all together, these results provide a comprehensive validation of the ROKS approach. We demonstrate that ROKS captures the essential physics of both the HAT and PCET photochemical pathways in liquid water. Based on the gas phase water dimer model, ROKS does not reproduce the conical intersection geometry for the HAT mechanism. In particular, a difference of approximately of $0.8$ Å is observed in the $\mathrm{HO}^{\bullet}$-$\mathrm{H}^{\bullet}$ distance at the conical intersection when compared with CASPT2. Nevertheless, ROKS qualitatively describes the same photodissociation pathway, reproducing both the slope of the excited-state potential energy surface and the presence of the relevant chemical species along the dissociation coordinate. While such differences may lead to deviations in the predicted HAT lifetimes in vacuum, we expect their impact to be reduced in the condensed phase. Indeed, in bulk liquid water, ROKS yields lifetimes and quantum yields that are in good agreement with experimental observations. Overall, the extensive validation presented here demonstrates the robustness of our methodology and reinforces confidence in its predictive capability for modeling the excited-state dynamics and photochemistry of liquid water. We nonetheless caution that more systematic studies with multireference methods for condensed phase aqueous systems needs to be explored in the future.